### v1

@@ 2.4.1 Latent allocations $\{z_{n}\}$
(kurihara2007collapsed). So, the variational distribution updates as
@@
where $N_{k}=\sum_{n}\mathbf{1}[z_{n}=k]$, $N_{>k}=\sum_{n}\mathbf{1}[z_{n}>k]$ and $N_{\geq k}=N_{k}+N_{>k}$. The calculations of their corresponding expectations and variances follow from kurihara2007collapsed (details in Supplementary material).
@@ 2.4.2 Distributional parameters $\{\mu_{k},\Sigma_{k}\}$
The updated $q^{*}(\mu_{k})$ depends on $p(\mu_{k})$ as well as $p(X_{n}=x_{n}|z_{n})$, and hence the structure of $\Sigma_{k}$ plays an important role. For instance, when we assume $\Sigma_{k}$ to be fixed and consider $\Sigma_{k}=\Sigma=\frac{1}{\sigma}I_{d\times d}$, the variational update of $\mu_{k}$ follows blei2006variational. On the other hand, if we consider $\Sigma_{k}=\Sigma$ as a full matrix, the updated hyper-parameters of $\mu_{k}$ can be written as
@@ 2.5 Implementation
Starting with an initial choice of latent allocation probabilities $q^{\circ}_{nk}$, the DPMM iteratively updates VI parameters based on the coordinate ascent algorithm (Eq: (2)) until there is no significant change in ELBO (blei2006variational)
@@
ELBO acts as a lower bound of log evidence $logp(X)$ which minimizes reverse KL divergence between the target and variational distributions. The objective of VI is to update the variational parameters such that ELBO is maximized, thus approximating $p(\cdot)$ with $q(\cdot)$.
Due to non-convexity of the ELBO, VI is sensitive to the choice of initial parameters (blei2017variational). We circumvent this by choosing random initialization, and select the optimum starting values based on variational log likelihood (VLL) term $\mathbb{E}_{q}[\log\,p(X|Z,\{\mu_{k},\Sigma_{k}\},\alpha)](=\mathbb{E}_{q}[\log\,p(X|Z)])$ from Eq: (10). Although ELBO explicitly regularizes model complexity by including negative entropy of $q(\cdot)$ via $\mathbb{E}_{q}[\log\,q(Z,\{\mu_{k},\Sigma_{k}\},\alpha)]$, it is effected numerically especially for higher dimensions and can over-influence the ELBO values (details in Supplementary material). The VLL, on the other hand, implicitly handles model complexity through expectation over $q(\cdot)$, while providing a numerical fit of the data based on updated variational distribution. Thus, it makes sense to select the initial VI parameters that maximizes VLL.
@@ 3 Results
The proposed DPMM integrates a hierarchical approach in terms of varying the DP concentration parameter $\alpha$ in a collapsed VI method. The variational Gamma hyper-parameter $w_{1}$ depends on the number of non-zero clusters $t$ present in Z (9a), instead of total variational clusters $K$. For a Mean field VI (blei2006variational), the distribution of Z depends on $\alpha$ via the mixing proportions $\{\pi_{k}\}$’s, which can be non-zero for empty clusters (smaller proportions for unlikely clusters). Owing to the hierarchical dependency between $\alpha$ and $\{\pi_{k}\}$’s, the variational update of $\alpha$ thus considers all the plausible clusters based on non-zero $\{\pi_{k}\}$’s (which is often equal to $K$), rather than only the non-zero clusters $t$. This leads to an inconsistency between the number of clusters estimated between $\alpha$ and $Z$ (details in Supplementary material). However, for a collapsed VI, the distribution of Z directly depends on $\alpha$ (1). And with the theoretical approximations (2.4.3), our approach is able to maintain consistent estimates of cluster number through the variational updates of $\alpha$ hyper-parameters (9a), (9b).
The cluster estimates also depend on the structure as well as prior choice of $G_{0}$ variance (hejblum2019sequential). A common covariance structure (models $M_{3},M_{4}$ and $M_{5}$ from Table 1) can prove inadequate to recover clusters with heterogeneous covariances (biernacki2002assessing). Alternatively, a cluster-specific covariance structure (models $M_{6},M_{7}$ and $M_{8}$) is favoured, which integrates within-cluster variance in the model and complements the effect of $\alpha$ (Fig: 1). For cluster-specific covariance structure, an Inverse-Wishart prior often incorporates rigidity in the model ($M_{6}$) (o2004kendall; tokuda2025visualizing). We introduced independence between variance and covariance terms by assigning element-wise prior distributions (models $M_{7}$ and $M_{8}$) and applied conjugate $\Gamma(.)$ distributions for the diagonal elements. For the off-diagonal elements, we considered Gaussian distributions with arbitrarily small fixed variance (model $M_{8}$) and 0-centered Laplace distribution (model $M_{7}$).
@@ 3.1 Effect of $G_{0}$ Covariance Structure and Prior Choice
In Figure 1, the fixed variance models $M_{1}$ with diagonal and $M_{2}$ with full covariance matrices, respectively, act as known parameters and set the baseline for performance comparison. For unknown variance, global covariance structure shows comparative performance with diagonal covariance model $M3$ and significantly poor performance with both full covariance model $M4$ and Cholesky-decomposed covariance model $M5$. $M3$ employs uncorrelated mixture distribution for all the variational clusters, thus reducing rigidity and variance-covariance coupling. Compared to global covariance, cluster-specific covariance structure provides better results.
@@
Element-wise prior models show better adaptability with varying cluster number than $IW$ prior model in a high-dimensional setting (Figure 2). $M_{7}$ and $M_{8}$ have comparable performances (with $M_{7}$ slightly better than $M_{8}$). This can be attributed to the fact that a Laplace distribution is equivalent to a marginalized Gaussian distribution with Exponential prior on the variance parameter. Owing to its sparsity inducing effect (jing2024variance), we call $M_{7}$ as the Sparse DPMM, and consider it to be the best possible choice among the models defined in Table 1.
3.2 Effect of hyper-parameters of cluster distributions $\{\mu_{k},\Sigma_{k}\}$
Real high-dimensional data suffers from the curse of dimensionality due to low signal-to-noise ratio and limited sample size. Sparse DPMM is equiped with hyper-parameters of $\{\mu_{k},\Sigma_{k}\}$ that can weed out features with weaker contributions, as well as increase the effective sample size, thus reducing model variability which is otherwise an issue for low-n-high-d settings. Due to standard procedure of normalising such data sets (xia2023statistical), we chose $\mu_{0}=\mathbf{0}$ as the prior hyper-parameter for $\mu_{k}$ and $a_{0}=b_{0}$ for the diagonal elements of $\Sigma_{k}^{-1}$ (2.4.2). Along with $a_{0}=b_{0}$, the scaling factor $k_{0}$ for $p(\mu_{k}|\Sigma_{k})\sim MVN(\mathbf{0},\frac{1}{k_{0}}\Sigma_{k}^{-1})$ (2.4.2) plays a significant role in defining the posterior cluster distributions via their variational updates (7), thus effecting posterior estimation of $K$.
We performed the sensitivity analysis with $N=100,d=1000$ Negative Binomial variables clustered into $K_{true}=3$ groups to illustrate the robustness of Sparse DPMM on non-Gaussian data with more than $2$ clusters. For $1000$ random initialisations, we compared the VLL values obtained for $a)$ different $a_{0}=b_{0}$ with fixed $k_{0}=1$ (Figure 3(a)) and $b)$ different $k_{0}$ with fixed $a_{0}=b_{0}=0.001$ (Figure 3(b)).
Figure 3: Sensitivity analysis of Sparse DPMM over the hyper-parameters 3(a) $a_{0}=b_{0}$ and 3(b) $k_{0}$ for Negative Binomial simulations with $K_{true}=3$
For Sparse DPMM, the derivatives of VLL with respect to $a_{0}=b_{0}$ and $k_{0}$ can be expressed as
$\frac{\delta}{\delta a_{0}}VLL$ $=\frac{1}{2}\sum_{k}\left(d\sum_{n}q_{nk}^{\circ}\Psi^{1}(a_{0}+\sum_{n}q_{nk}^{\circ})\right$.
$\left.-\sum_{i}\biggl[\frac{a_{0}\sum_{n}q_{nk}^{\circ}+(\frac{\sum_{n}q_{nk}^{\circ}x_{n_{i}}^{2}}{2})^{2}-\frac{\sum_{n}q_{nk}^{\circ}\sum_{n}q_{nk}^{\circ}x_{n_{i}}^{2}}{2}}{(a_{0}+\frac{1}{2}\sum_{n}q_{nk}^{\circ}x_{n_{i}}^{2})^{2}}\biggr]\right)$ (11a)
$\frac{\delta}{\delta k_{0}}VLL$ $=\sum_{k}\biggl[\frac{1}{(k_{0}+\sum_{n}q_{nk}^{\circ})^{2}}\left(d\sum_{n}q_{nk}^{\circ}-(2-\sum_{n}q_{nk}^{\circ})\sum_{i}(\sum_{n}q_{nk}^{\circ}x_{n_{i}})^{2}\right)\biggr]$ (11b)
where $\{q_{nk}^{\circ}\}$ are a random allocation probability values and $\Psi^{1}(\cdot)$ is the trigamma function (details in Supplementary material). Functional behaviour of $\frac{\delta}{\delta a_{0}}VLL$ relies on $\{q_{nk}^{\circ}\}$, and VLL can either attain maxima or converge to a stabilising value with increasing $a_{0}=b_{0}$ (Figure 3(a)). For both the conditions, we estimate $a_{0}$ for which either $\frac{\delta}{\delta a_{0}}VLL=0$ or the maximum curvature point of $\frac{\delta}{\delta a_{0}}VLL$ that corresponds to the stabilising convergence of VLL (details in Supplementary material). In either case, the estimated $a_{0}$ represents the optimal value after which there is no further gain in the log likelihood, hence no additional prior information can be incorporated with higher $a_{0}$. Thus, for a random choice of $\{q_{nk}^{\circ}\}$, $a_{0}$ estimation inherently puts an upper bound on the degree of information required for maximising the VLL. This enables the model to threshold dimensions that have stronger contributions to effect the posterior estimation, and subsequently the VLL for a given data.
On the other hand, behaviour of $\frac{\delta}{\delta k_{0}}VLL$ is converging, and the value of convergence depends on $\{q_{nk}^{\circ}\}$ (Figure 3(b)). $k_{0}$ acts as the equivalent sample size of the $\mu_{k}$ prior (murphy2007conjugate), and scales the variational hyper-parameters of $\mu_{k}$ in Eq: (6). So, a smaller value of $\frac{1}{k_{0}}$ reduces the variability of $\mu_{k}$ which is equivalent of having a larger data sample. As the variational updates of $\mu_{k}$ in Eq: (6) is a weighted average between cluster-specific samples and prior mean, a natural choice is to consider $k0=N+1$ where N is the data sample size.
@@ 3.3 Computational scalability
We implemented our model framework in an R-package vimixr with all the $8$ parameterisations given in Table 1. For Sparse DPMM, the computation scales with sample size $N$ as order $\mathcal{O}(Nlog(N))$ due to the allocation variable $\mathbf{Z}$ updates with conditional expectation on $z_{n}|\mathbf{z}^{-n}$. The computational dependence for dimension $d$ is of order $\mathcal{O}(dlog(d)log(log(d))$ (details in Supplementary material). Although higher dimensions achieve convergence with less number of iterations, the rate is slightly higher due to computation of cluster-specific $d$-dimensional precision matrices.
For a slice-sampling based MCMC approach, we used the function DPMGibbsN from NPflow R package (hejblum2019sequential) and compared our Sparse DPMM for $N=100,d=100$ Gaussian data. Even with $1000$ MCMC iterations, DPMGibbsN often fails to achieve convergence. Sparse DPMM, on the other hand, demonstrates an increase in computational time by a magnitude of two orders ($\sim 100$ times) during bench-marking (details in Supplementary material).
@@ 3.4 Cancer sub-type estimation
We applied the proposed clustering framework on the gene-expression data set from leukemia sub-type study by armstrong2002mll, previously used as a benchmark for comparative analysis of clustering algorithms de2008clustering. armstrong2002mll showed that leukemia with rearrangement of the MLL gene (mixed lineage leukemia gene, now renamed KMT2A) should be classified as a distinct clinical entity rather than common acute lymphoblastic leukemia. The chromosomal rearrangement of MLL creates fusion genes with various partner genes that influence lineage commitment (krivtsov2007mll). For example, certain MLL rearrangements result in overexpression of HOXA9 and PRG1 genes, which are characteristic markers of acute myeloid leukemia (AML) rather than ALL (armstrong2002mll). Recent studies have further revealed the lineage plasticity of MLL-rearranged leukemia, showing dynamic transitions between lymphoid and myeloid phenotypes (chen2022single; janssens2024mll).
Working on the Affymetrix data from armstrong2002mll, de2008clustering provides a filtered set of $d=2194$ expressed genes for clustering $N=72$ leukemia samples into ALL-MLL-AML sub-types (Fig 4(a)). For a strong choice of prior hyper-parameters corresponding to Section 3.2, we obtained $K_{post}=3$ with ARI score of $0.92$ (Figure 4(b)).
@@
Following our empirical Bayes approach in Section 3.2, we estimated $a_{0}=b_{0}=28.9076$, and along with $k_{0}=N+1=73$, we applied Sparse DPMM for random initial probability allocation values and chose the best performing model based on VLL. To check the sensitivity of Sparse DPMM with a weaker prior hyper-parameter, we used $a_{0}=b_{0}=10$ instead of $50$ and clustered the same data set. For this particular choice of $a_{0}=b_{0}$, we obtained $K_{post}=4$ with ARI score $0.86$. (Figure 5).
Figure 5: PCA projection on the first 22 principal components for Sparse DPMM cluster estimates obtained using weaker hyper-prarameters
Figure 6: Heatmap showing comparison between labelled sub-types and Sparse DPMM (weaker) cluster estimates with respect to specifically expressed genes for the labelled sub-types (armstrong2002mll) for all the $72$ samples (genes in bold, probe-id in italics)
Comparing with known sub-types in Figure 4(a), Sparse DPMM with weaker hyper-parameters identifies an additional $4^{th}$ cluster with one sample replacing a labelled AML sub-type (Figure 5). Analogous to (armstrong2002mll), we plot a heat-map based on the available genes (de2008clustering) that are specifically expressed for the known sub-types to compare the lineage of the $4^{th}$ estimated cluster (Figure 6). The heat-map reveals a mixed expression pattern for the 4th cluster. The cluster shows elevated expression of genes characteristic of ALL sub-types (CD22 and CD24), suggesting lymphoblastic features. It also displays moderate expression of genes typically overexpressed in MLL sub-types (PROML1, FLT3 and ADAM10). In contrast, genes predominantly expressed in AML sub-types (DF, CTSD and BSG) are not dominantly expressed in this cluster. However, this expression pattern is not clear-cut, with the 4th cluster displaying characteristics intermediate between ALL and MLL rather than fitting cleanly into a single category. This ambiguous profile is consistent with the documented lineage plasticity of MLL-rearranged leukemias, which can undergo dynamic transitions between lymphoid and myeloid phenotypes during disease progression and treatment (chen2022single; janssens2024mll). The identification of this 4th cluster by Sparse DPMM thus appears to capture this biological reality of MLL heterogeneity rather than representing a technical artifact.

### v2

@@ 2.4.1 Latent allocations $\{z_{n}\}$
(Kurihara et al., 2007). So, the variational distribution updates as
@@
where $N_{k}=\sum_{n}\mathbf{1}[z_{n}=k]$, $N_{>k}=\sum_{n}\mathbf{1}[z_{n}>k]$ and $N_{\geq k}=N_{k}+N_{>k}$. The calculations of their corresponding expectations and variances follow from Kurihara et al. (2007) (details in Supplementary material).
@@ 2.4.2 Distributional parameters $\{\mu_{k},\Sigma_{k}\}$
The updated $q^{*}(\mu_{k})$ depends on $p(\mu_{k})$ as well as $p(X_{n}=x_{n}|z_{n})$, and hence the structure of $\Sigma_{k}$ plays an important role. For instance, when we assume $\Sigma_{k}$ to be fixed and consider $\Sigma_{k}=\Sigma=\frac{1}{\sigma}I_{d\times d}$, the variational update of $\mu_{k}$ follows Blei and Jordan (2006). On the other hand, if we consider $\Sigma_{k}=\Sigma$ as a full matrix, the updated hyper-parameters of $\mu_{k}$ can be written as
@@ 2.5 Implementation
Starting with an initial choice of latent allocation probabilities $q^{\circ}_{nk}$, the DPMM iteratively updates VI parameters based on the coordinate ascent algorithm (Eq: (2)) until there is no significant change in ELBO (Blei and Jordan, 2006)
@@
ELBO acts as a lower bound of log evidence $logp(X)$ which minimizes reverse KL divergence between the target and variational distributions. Hence, objective of VI is to update the variational parameters such that ELBO is maximized, thus approximating $p(\cdot)$ with $q(\cdot)$.
Due to non-convexity of the ELBO, VI is sensitive to the choice of initial parameters (Blei et al., 2017). We circumvent this by choosing random initialization, and select the optimum starting values based on variational log likelihood (VLL) term $\mathbb{E}_{q}[\log\,p(X|Z,\{\mu_{k},\Sigma_{k}\},\alpha)](=\mathbb{E}_{q}[\log\,p(X|Z)])$ from Eq: (10). Although ELBO explicitly regularizes model complexity by including negative entropy of $q(\cdot)$ via $\mathbb{E}_{q}[\log\,q(Z,\{\mu_{k},\Sigma_{k}\},\alpha)]$, it can over-influence the ELBO due to large number of parameters in high dimensions (details in Supplementary material). The VLL, on the other hand, implicitly handles model complexity through expectation over $q(\cdot)$, while providing a numerical fit of the data based on updated variational distribution.
@@ 3 Results
The proposed DPMM integrates a hierarchical approach in terms of varying the DP concentration parameter $\alpha$ in a collapsed VI method. The variational Gamma hyper-parameter $w_{1}$ depends on the number of non-zero clusters $t$ present in Z (9a), instead of total variational clusters $K$. For a Mean field VI (Blei and Jordan, 2006), the distribution of Z depends on $\alpha$ via the mixing proportions $\{\pi_{k}\}$’s, which can be non-zero for empty clusters (smaller proportions for unlikely clusters). Owing to the hierarchical dependency between $\alpha$ and $\{\pi_{k}\}$’s, the variational update of $\alpha$ thus considers all the plausible clusters based on non-zero $\{\pi_{k}\}$’s (which is often equal to $K$), rather than only the non-zero clusters $t$. This leads to an inconsistency between the number of clusters estimated between $\alpha$ and $Z$ (details in Supplementary material). However, for a collapsed VI, the distribution of Z directly depends on $\alpha$ (1). And with the theoretical approximations (2.4.3), our approach is able to maintain consistent estimates of $K$ through the variational updates of $\alpha$ hyper-parameters (9a), (9b).
The cluster estimates also depend on the structure as well as prior choice of $G_{0}$ variance (Hejblum et al., 2019). A common covariance structure (models $M_{3},M_{4}$ and $M_{5}$ from Table 1) can prove inadequate to recover clusters with heterogeneous covariances (Biernacki et al., 2002). Alternatively, a cluster-specific covariance structure (models $M_{6},M_{7}$ and $M_{8}$) is favoured, which integrates within-cluster variance in the model and complements the effect of $\alpha$ (Fig: 1). For cluster-specific covariance structure, an Inverse-Wishart prior often incorporates rigidity in the model ($M_{6}$) (Tokuda et al., 2025). We introduced independence between variance and covariance terms by assigning element-wise prior distributions (models $M_{7}$ and $M_{8}$) and applied conjugate $\Gamma(.)$ distributions for the diagonal elements. For the off-diagonal elements, we considered Gaussian distributions with arbitrarily small fixed variance (model $M_{8}$) and 0-centered Laplace distribution (model $M_{7}$).
@@ 3.1 Effect of $G_{0}$ Covariance Structure and Prior Choice
In Figure 1, the fixed variance models $M_{1}$ with diagonal and $M_{2}$ with full covariance matrices act as known parameters and set the baseline for performance comparison. For unknown variance, global covariance structure shows comparative performance with diagonal covariance $M3$ and significantly poor performance with both full covariance $M4$ and Cholesky-decomposed covariance $M5$. $M3$ employs uncorrelated mixture distribution for all the variational clusters, thus reducing rigidity and variance-covariance coupling. Compared to global covariance, cluster-specific covariance structure provides better results.
@@
Element-wise prior models show better adaptability with varying cluster number than $IW$ prior model in a high-dimensional setting (Figure 2). $M_{7}$ and $M_{8}$ have comparable performances (with $M_{7}$ slightly better than $M_{8}$). This can be attributed to the fact that a Laplace distribution is equivalent to a marginalized Gaussian distribution with Exponential prior on the variance parameter. Owing to its sparsity inducing effect (Jing et al., 2024), we call $M_{7}$ as the Sparse DPMM, and consider it to be the best possible choice among the models defined in Table 1.
Analogous to real-world datasets, we sampled Negative Binomial simulations to study the effect of $G_{0}$ hyper-parameters for Sparse DPMM. Analytical calculations provide data-dependent updates, which play the role of empirical Bayes estimators for those hyper-parameters (details in Supplementary materials).
@@ 3.2 Computational scalability
We implemented our model framework in an R-package vimixr with all the $8$ parameterizations given in Table 1. For Sparse DPMM, the computation scales with sample size $N$ as order $\mathcal{O}(Nlog(N))$ due to the allocation variable $\mathbf{Z}$ updates with conditional expectation on $z_{n}|\mathbf{z}^{-n}$. The computational dependence for dimension $d$ is of order $\mathcal{O}(dlog(d)log(log(d))$ (details in Supplementary material). Although higher dimensions achieve convergence with less number of iterations, the rate is slightly higher due to computation of cluster-specific $d$-dimensional precision matrices.
For a slice-sampling based MCMC approach, we used the function DPMGibbsN from NPflow R package (Hejblum et al., 2019) and compared our Sparse DPMM for $N=100,d=100$ Gaussian data. Even with $1000$ MCMC iterations, DPMGibbsN often fails to achieve convergence. Sparse DPMM, on the other hand, demonstrates an increase in computational time by a magnitude of two orders ($\sim 100$ times; details in Supplementary material).
@@ 3.3 Cancer sub-type estimation
We applied the proposed clustering framework on the gene-expression dataset from leukemia sub-type study by Armstrong et al. (2002), previously used as a benchmark for comparative analysis of clustering algorithms (De Souto et al., 2008). Armstrong et al. (2002) showed that leukemia with rearrangement of the MLL gene (mixed lineage leukemia gene, now renamed KMT2A) should be classified as a distinct clinical entity rather than common acute lymphoblastic leukemia. The chromosomal rearrangement of MLL creates fusion genes with various partner genes that influence lineage commitment (Krivtsov and Armstrong, 2007). For example, certain MLL rearrangements result in overexpression of HOXA9 and PRG1 genes, which are characteristic markers of acute myeloid leukemia (AML) rather than ALL (Armstrong et al., 2002). Recent studies have further revealed the lineage plasticity of MLL-rearranged leukemia, showing dynamic transitions between lymphoid and myeloid phenotypes (Chen et al., 2022; Janssens et al., 2024).
Working on the Affymetrix data from Armstrong et al. (2002), De Souto et al. (2008) provides a filtered set of $d=2194$ expressed genes for clustering $N=72$ leukemia samples into ALL-MLL-AML sub-types (Fig 3(a)). For a strong choice of prior hyper-parameters, we obtained $K_{post}=3$ with ARI score of $0.92$ (Figure 3(b)).
@@
Following our empirical Bayes approach, we estimated $a_{0}=b_{0}=28.9076$, and along with $k_{0}=N+1=73$, we applied Sparse DPMM for random initial probability allocation values and chose the best performing model based on VLL. To check the sensitivity of Sparse DPMM with a weaker prior hyper-parameter, we used $a_{0}=b_{0}=10$ instead of $50$ and clustered the same dataset. For this particular choice of $a_{0}=b_{0}$, we obtained $K_{post}=4$ with ARI score $0.86$. (Figure 4(a)).
(a) Sparse DPMM clusters using weaker hyper-prarameters
(b) Heat map of clusters based on sub-type specific expressed genes
Figure 4: 4(a)) PCA projection on the first 22 principal components for Sparse DPMM clusters obtained using weaker hyper-prarameters; 4(b)) Heatmap showing comparison between labelled sub-types and Sparse DPMM (weaker) cluster estimates with respect to specifically expressed genes for the labelled sub-types (Armstrong et al., 2002) for all the $72$ samples (genes in bold, probe-id in italics)
Comparing with known sub-types in Figure 3(a), Sparse DPMM with weaker hyper-parameters identifies an additional $4^{th}$ cluster with one sample replacing a labelled AML sub-type (Figure 4(a)). Analogous to (Armstrong et al., 2002), we plot a heat-map based on the available genes (De Souto et al., 2008) that are specifically expressed for the known sub-types to compare the lineage of the $4^{th}$ estimated cluster (Figure 4(b)). The heat-map reveals a mixed expression pattern for the 4th cluster. The cluster shows elevated expression of genes characteristic of ALL sub-types (CD22 and CD24), suggesting lymphoblastic features. It also displays moderate expression of genes typically overexpressed in MLL sub-types (PROML1, FLT3 and ADAM10). In contrast, genes predominantly expressed in AML sub-types (DF, CTSD and BSG) are not dominantly expressed in this cluster. However, this expression pattern is not clear-cut, with the 4th cluster displaying characteristics intermediate between ALL and MLL rather than fitting cleanly into a single category. This ambiguous profile is consistent with the documented lineage plasticity of MLL-rearranged leukemias, which can undergo dynamic transitions between lymphoid and myeloid phenotypes during disease progression and treatment (Chen et al., 2022; Janssens et al., 2024). The identification of this 4th cluster by Sparse DPMM thus appears to capture this biological reality of MLL heterogeneity rather than representing a technical artifact.